Practice: Determine the molarity of a solution formed by dissolving 468 mg of MgI 2 in enough water to yield 50.0 mL of solution.
Answer #1

Mass of MgI2 = 468 mg × (0.001g/1mg) = 0.468 g

Moles of MgI2 = mass(g)/molar mass(g/mol) = 0.468g/278.114g/mol = 0.001683 moles MgI2(solute) .

Volume of solution = 50mL × (0.001L/1mL) = 0.050L

Molarity of solution = moles of solute/volume of solution in litres = 0.001683moles/0.050L = 0.0337 M

Thus molarity of solution = 0.0337 M .

Kindly rate the answer.